# How Well do LLMs Compress Their Own Chain-of-Thought? A Token Complexity Approach

### v1

@@ 4 Theoretical Limits of the Length-Performance Tradeoff
Not only does token-complexity provide an interpretable and accurate model of reasoning task performance, it gives insight into how to improve the token-length efficiency. Viewing these response-reducing CoT prompts as a form of lossy compression, this raises a natural research question: what are the theoretical limits on the optimal tradeoff token-length and accuracy and how far away are existing prompting strategies from this limit? Although there are recent works which propose methodologies for improving this tradeoff, so far there has not been a study of the theoretical limits of compression, which makes it difficult to evaluate these proposed methods.
We develop a framework to empirically compute information-theoretic lower bounds on performance, inspired by rate-distortion theory. First, we define $\bar{t}_{\pi}(P)\equiv\frac{1}{n}\sum_{i=1}^{n}t(X_{i})$ to be the average token-length under CoT prompt $P$ and LLM $\pi$. We define $\alpha^{*}(T)$ to be the optimal performance for an average token budget of $T$ tokens, and $T^{*}(\alpha)$ is the minimum average token length to achieve a target accuracy of $\alpha$:
@@
These quantities involve intractable optimizations over CoT prompts $P$. Yet, under the token complexity hypothesis, these optimization problems are greatly simplified, involving only optimization over token-lengths instead of prompts.
In fact, the optimization is a special case of a knapsack problem, and thus gives rise to a closed-form solution. First, define the empirical CDF of the true token complexities $F_{n}(t)\equiv\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\tau_{i}^{\pi}\leq t\}$, let $E_{n}(t)\equiv\frac{1}{n}\sum_{i=1}^{n}\tau_{i}^{\pi}\mathbf{1}\{\tau_{i}^{\pi}\leq t\}$, and let $\bar{\tau}_{\pi}=\frac{1}{N}\sum_{i=1}^{n}\tau_{i}^{\pi}\mathbf{1}\{\tau_{i}^{\pi}<\infty\}$ denote the average token complexity among questions that have finite token complexity.
@@
Figure 3: Do LLM’s tailor response length to problem difficulty? (Left) Average token length produced by GPT-4o on MMLU-Pro Math across prompts in Table 1, split by problems which can be solved without chain-of-thought and problems which NoCoT does not successfully solve. Response lengths are consistently higher for the problems NoCoT unsuccessfully solves. (Right) Token lengths across 500 questions in MMLU-Pro Math for GPT-4o under the DefaultCoT, BeConcise, and WordLimit(10) prompts. Questions are sorted in increasing order of token-complexity across different prompt IDs. Token lengths exhibit a low but non-zero correlation with the true token complexity.
@@
Nevertheless $\alpha^{*}_{\pi}(T)$ provides a computable upper bound on maximum accuracy for a particular token budget, which plot in Figure 1 under the label ‘oracle upper bound’, using estimated token complexities $\hat{\tau}_{i}^{\pi}$. Across all the LLMs and benchmarks we consider, we find that this is indeed serves as an upper bound on the performance of the prompting strategies we consider across response lengths, especially for more difficult datasets such as MATH-500 and MMLU-Pro Math. Yet, for GSM8K in Appendix E, we see that this gap with the upper bound $\alpha^{*}_{\pi}(T)$ is much smaller, illustrating that while $\alpha^{*}_{\pi}(T)$ may be challenging to achieve it still gives a reasonable upper bound on performance.
Finally, we circle back to the observation made in Section 2.2 that one can substantially reduce the length of the chain-of-thought (with prompts like BeConcise or NoProperGrammar) while maintaining a similar accuracy to DefaultCoT. This leads to a natural question: what is the lower bound on the number of tokens needed in order to achieve the best possible accuracy? Under the token-complexity hypothesis, we can characterize a simple closed-form expression for this lower bound.
@@
Corollary 1 gives the number of tokens required for ‘lossless compression’, i.e. achieving the best possible accuracy. Remarkably, this lower bound is only the mean of the token-complexities, which can be much smaller than the average token-length of existing prompting strategies. In Table 3, we consider the MMLU-Pro Math dataset across several LLMs, and we compare the lower bound $T_{\pi}^{*}(A^{*})$ with the token counts of DefaultCoT and BeConcise. Across all the LLMs, BeConcise reduces the token counts by 1.2-1.4x, while we see that the theoretical upper bound on token reduction ranges from 3.2-11.2x. Along with the results in Figure 1, this illustrates that while the upper bound may not be exactly attainable, there may be significant room for improvement.
5 Evaluation of Adaptive Compression
Most of all, the token complexity hypothesis emphasizes the importance of adaptive compression: under Assumption 1, improvements in compression only come from matching the token-length $t(X_{i})$ to the token complexity $\tau_{i}^{\pi}$ as closely as possible, rather than from modifications in the chain-of-thought $X_{i}$ itself.
We proceed to show how token complexity provides a framework for studying the capabilities of existing language models to tailor their response length to problem difficulty. Our main observation is that across the chains-of-thought we consider, LLMs do natively adjust response length to the difficulty of the problem, without explicitly being asked to. However, this adaptivity is limited and far from optimal.
Our first evaluation of this capability compares performance of our prompts with a non-adaptive baseline: how does the accuracy-token length trade-off curve compare with a non-adaptive strategy that uses exactly the same number of tokens per question? Note that simply prompting the LLM to use a fixed number of tokens/words/characters/etc., is unable to force a fixed response length Han et al. (2024). To construct this baseline, we let $\bar{\alpha}_{\pi}(T)=\frac{1}{n}\sum\nolimits_{i=1}^{n}\mathbf{1}\{\tau^{\pi}_{i}\leq T\}$ denote the accuracy under the token complexity hypothesis if we use exactly $T$ tokens for each question. Since these values describe the performance under a sub-optimal, non-adaptive strategy, $\bar{\alpha}_{\pi}(T)\leq\alpha^{*}(T)$. The blue line in Figure 1 displays $\bar{\alpha}_{\pi}(T)$ across average token lengths $T$. We observe that across LLMs and benchmarks, the performance of the prompts is close to the non-adaptive baseline $\bar{\alpha}_{\pi}(T)$, with some prompts such as WordLimit(k) or OnlyNumbers exceeding $\bar{\alpha}_{\pi}(T)$. This provides evidence that LLMs outperform a non-adaptive baseline, but not by much.
Model MMLU-Pro Math GSM8K MATH-500 GPT-4o 92.1% 97.0% 89.9% GPT-4o Mini 91.3% 94.1% 91.4% Claude 3.5 Sonnet 92.6% 97.4% 92.1% Claude 3.5 Haiku 90.1% 94.8% 90.2% Llama 3.3 70B 90.6% 96.2% 89.8%
Table 4: Average classification accuracy $\bar{c}_{\pi}$ of threshold classifier $\mathbf{1}\{t(X_{i,k})>\hat{\tau}_{i}^{\pi}\}$ across LLMs and reasoning benchmarks.
Model MMLU-Pro Math GSM8K MATH-500 GPT-4o 5.0% 1.6% 6.0% GPT-4o Mini 5.2% 5.1% 5.7% Claude 3.5 Sonnet 3.4% 1.4% 4.5% Claude 3.5 Haiku 5.1% 4.1% 6.3% Llama 3.3 70B 5.1% 2.7% 6.1%
Table 5: Average relative discrepancy $\text{Err}_{\pi}$ across LLMs and benchmarks. $\text{Err}_{\pi}$ is relative absolute loss between actual performance $\text{Acc}_{\pi}(P_{k})$ and predicted performance $\widehat{\text{Acc}}_{\pi}(P_{k})$, averaged across prompts $P_{k}$.
We can study this in more granularity at the question-level. On the left panel of Figure 3, we consider the average token lengths of GPT-4o on MMLU-Pro Math. We split the problems into two categories: problems that the LLM successfully solves without chain-of-thought (NoCot), and the problems NoCot unsuccessfully solves. Intuitively, the first class of problems are ‘easy’, since they do not require any chain-of-thought, and the rest are harder. Across all the prompts we consider, we observe that the average token-length among ‘easy’ problems is universally smaller than among ‘harder’ problems. Surprisingly, this is even true for the WordLimit/TokenLimit/CharLimit/etc. prompts, which are supposed to limit the LLM’s response to a fixed length. Nevertheless, despite the evidence of adaptive response lengths, this also shows that there is a lot of room for improvement: even though the LLM can solve the problem without any chain-of-thought, the model still proceeds to generate long chains-of-thought.
Through our estimated token complexities, we can perform a much more fine-grained analysis of adaptive response lengths. In the right panel of Figure 3, we look at the token-lengths across questions in MMLU-Pro Math induced by several CoT Prompts – DefaultCoT, BeConcise, and WordLimit(10) – for GPT-4o and Claude 3.5 Sonnet. We sort the questions in increasing order of the estimated token complexities $\tau_{i}^{\pi}$ to assess the degree to which the token lengths produced by the LLM correlate with the true token complexities. We can confirm the observation made earlier: the token lengths exhibit a low but non-zero correlation with the true token complexity. Table 6 in Appendix A measures Kendall-$\tau$ correlation between the token complexities and token lengths across questions for GPT-4o on MMLU-Pro Math, with the highest correlation of 0.53 achieved by OnlyNumbers and the lowest correlation of 0.04 achieved by NoCoT. In total, we highlight that token complexity allows researchers to study the ability to adapt response length to the difficulty of the problem by providing a finer-grained metric of problem difficulty.
@@ 6 Towards the Theoretical Limit
The goal of this work is not to propose a new methodology for shortening the response length, and instead our main contribution is to (1) use token complexity to assess existing compression strategies and (2) establish theoretical upper bounds for the optimal tradeoff between reasoning length and performance. Nonetheless, we will briefly discuss how improving adaptivity through predicting problem difficulty can improve upon the universal tradeoff curve observed in in Figure 1 and approach the upper bound $\alpha_{\pi}^{*}(T)$.
Figure 4: Average token length vs Accuracy for Claude 3.5 Sonnet on MMLU-Pro-Math under different prompt routing strategies. Unvertainty-based routing (in purple) does not improve much over existing prompting strategies. However Verifier-based routing (in blue) is able to achieve a better Pareto-frontier, closer to the theoretical upper bound.
As a proof of concept, we consider two prompt-routing strategies, in a similar spirit to Han et al. (2024). First, following the literature on calibration of LLMs Kadavath et al. (2022); Zhu et al. (2023); Chen et al. (2023), we use the language model probabilities in order to use a longer response if the model is uncertain about its current answer. Specifically, the LLM is first prompted with NoCot, i.e. to answer the question without thinking. Then we query the LLM’s token probability of their answer. If the probability is strictly below a threshold, we then query the LLM again with CoT prompt designed to get a longer response (e.g. DefaultCoT). The purple dots in Figure 1 show the performance of this prompt-routing scheme for Claude 3.5 Sonnet on MMLU-Pro Math, where we set the longer prompt to be a selection of four prompts. We observe that the performance is incrementally better but not much so. Moreover, these results are under a very high threshold: we only stick with the answer from NoCoT only if the probability is equal to 1.0. This illustrates that while there are benefits for adapting response length, these benefits may be limited unless one is able to identify question difficulty with high precision.
Nonetheless, if one has access to a perfect verifier, this can substantially improve the compression tradeoff. To evaluate this, we consider another prompt routing strategy, where one instead can verify whether the answer produced by the LLM under NoCoT is correct or not. If the answer is incorrect, then we route to a longer prompt. The blue dots in Figure 4 show performance of verifier-based routing with the same four longer prompts. Unlike uncertainty-based routing, these prompting strategies achieve a significantly better trade-off between reasoning length and performance, and approach the theoretical upper bound. While access to a verifier is unrealistic in practice, this shows how better information of problem difficulty leads to substantial improvements in chain-of-thought compression. This also illustrates that the upper bound $\alpha_{\pi}^{*}(T)$ can be approached and does accurately signify the potential improvements in the accuracy-length tradeoff that can be achieved. Altogether, this demonstrates that our framework sets natural baselines for researchers to compare the performance of length-compressed schemes.
@@ 7 Conclusion
Our study presents a systematic investigation into the trade-off between reasoning length and performance in large language models (LLMs), across different prompts. We demonstrate that this trade-off follows a universal Pareto curve, suggesting that reasoning length, rather than specific compression strategies, primarily determines accuracy. Introducing the concept of token complexity, we find that LLM performance at the question-level exhibits a sharp threshold behavior – it is successful only if its chain-of-thought is above a token threshold. Our theoretical analysis, grounded in rate-distortion theory, reveals that existing prompt-based compression strategies operate far from the optimal accuracy-length frontier, highlighting substantial room for improvement. Our work enables researchers to contextualize the performance of new methodologies for improving chain-of-thought compression and assess adaptivity.
Limitations
@@
Appendix A Additional Tables
The following is a table describing the Kendall-$\tau$ correlation between token-lengths with token-complexity across different prompts for GPT-4o on MMLU-Pro Math. While the ordering of which prompts have the highest correlation changes across different models and benchmarks, we observe widely that the correlation ranges between 0 - 0.5.
@@ Appendix A Additional Tables
Kendall
Correlation
OnlyNumbers 0.53
NoSpaces 0.51
CharLimit(10) 0.50
CharLimit(50) 0.49
NoProperGrammar 0.48
WordLimit(15) 0.25
TokenLimit(10) 0.25
WordLimit(5) 0.20
WordLimit(10) 0.18
NoCoT 0.04
Table 6: GPT-4o on MMLU-Pro Math: Kendall-$\tau$ correlation of token-lengths with token-complexity across different prompts.
@@
Appendix B Proof of Theorem 1
@@
Appendix C Example Prompt
@@
Answer the following question. {PROMPT}
Question: {QUESTION}
The last line of your response
should be of the following format:
Answer: \$ANSWER (without quotes)
where ANSWER is your final answer.
Appendix D Tradeoff curves for more models and benchmarks

### v2

@@ 4 Theoretical Limits of the Length-Performance Tradeoff
Not only does token-complexity provide an interpretable and accurate model of reasoning task performance, it gives insight into how to improve efficiency. Viewing these response-reducing CoT prompts as a form of lossy compression, this raises a natural research question that has so far not been studied: what are the theoretical limits on the optimal tradeoff token-length and accuracy and how far away are existing prompting strategies from this limit?
We develop a framework to empirically compute bounds on compression performance, inspired by rate-distortion theory. First, we define $\bar{t}_{\pi}(P)\equiv\frac{1}{n}\sum_{i=1}^{n}t(X_{i})$ to be the average token-length under CoT prompt $P$ and LLM $\pi$. We define $\alpha^{*}(T)$ to be the optimal performance for an average token budget of $T$ tokens, and $T^{*}(\alpha)$ is the minimum average token length to achieve an accuracy of $\alpha$:
@@
These quantities involve intractable optimizations over CoT prompts $P$. Yet, under the token complexity hypothesis, these optimization problems are greatly simplified, involving only optimization over token-lengths instead of prompts. In fact, the optimization problem is a special case of a knapsack problem, and thus gives rise to a closed-form solution. First, define the empirical CDF of the true token complexities $F_{n}(t)\equiv\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\tau_{i}^{\pi}\leq t\}$, let $E_{n}(t)\equiv\frac{1}{n}\sum_{i=1}^{n}\tau_{i}^{\pi}\mathbf{1}\{\tau_{i}^{\pi}\leq t\}$, and let $\bar{\tau}_{\pi}=\frac{1}{N}\sum_{i=1}^{n}\tau_{i}^{\pi}\mathbf{1}\{\tau_{i}^{\pi}<\infty\}$ denote the average token complexity among questions that have finite token complexity.
@@
Nevertheless $\alpha^{*}_{\pi}(T)$ provides a computable upper bound on maximum accuracy for a particular token budget, which we plot in Figure 1 under the label ‘oracle upper bound’, using estimated token complexities $\hat{\tau}_{i}^{\pi}$. Across all the LLMs and benchmarks we consider, we find that this is indeed serves as an upper bound on the performance of the prompting strategies we consider across response lengths, especially for more difficult datasets such as MATH-500 and MMLU-Pro Math. Yet, for GSM8K in Appendix G, we see that this gap with the upper bound $\alpha^{*}_{\pi}(T)$ is much smaller, illustrating that while $\alpha^{*}_{\pi}(T)$ may be challenging to achieve it still gives a reasonable upper bound on performance.
Finally, we circle back to the observation made in Section 2.2 that one can substantially reduce the length of the chain-of-thought (with prompts like BeConcise or NoProperGrammar) while maintaining a similar accuracy to DefaultCoT. This leads to a natural question: what is the lower bound on the number of tokens needed in order to achieve the best possible accuracy? Under the token-complexity hypothesis, we obtain a simple closed-form expression for this lower bound.
@@
Corollary 1 gives the number of tokens required for ‘lossless compression’, i.e. achieving the best possible accuracy. Remarkably, this lower bound is only the mean of the token-complexities, which can be much smaller than the average token-length of existing prompting strategies. In Table 4, we consider the MMLU-Pro Math dataset across several LLMs, and we compare the lower bound $T_{\pi}^{*}(A^{*})$ with the token counts of DefaultCoT and BeConcise. Across all the LLMs, while BeConcise reduces the token counts by 1.2-1.4x, the token reduction achieved by the optimal compression scheme is 3.2-11.2x, and is especially high for GSM8K. Along with the results in Figure 1, this illustrates that while the upper bound may not be exactly attainable, there may be significant room for improvement especially for easier datasets.
Finally, we note that token complexity may be of broader interest as a measure of reasoning capabilities, even for benchmarks which are ‘saturated’. For instance, both Claude 3.5 Sonnet and Claude 3.5 Haiku achieve a similar accuracy on GMS8K (97% vs 95% under DefaultCoT). Yet the average token complexity $\bar{\tau}_{\pi}$ is 42.4 for Haiku and 17.9 for Sonnet, showing that even if the accuracy is similar, Sonnet can achieve it with much fewer tokens.
@@ 5 Towards the Theoretical Limit
The token complexity hypothesis not only provides limits on the efficiency of an optimal chain-of-thought compression scheme, it highlights the importance of adaptive compression – using shorter chains-of-thought for easier questions – for approaching these limits. This provides theoretical motivation for methods developed in recent works (e.g. Han et al. (2024); Kang et al. (2024)) designed to calibrate the length of the chain-of-thought to problem difficulty. In this section, we illustrate how our empirical framework can help evaluate and contextualize recent advances in adaptive reasoning.
Figure 3: Does adaptive compression improve the accuracy-length tradeoff? (Left) Average token length vs Accuracy for GPT-4o on MMLU-Pro-Math under different prompt routing strategies. TALE-EP (Han et al., 2024) (light green), which first has the LLM guess the number of tokens to use, is slightly below the Pareto curve achieved by simpler prompting strategies. Only verifier-based routing (in blue) is able to achieve a better tradeoff, closer to the upper bound. (Right) We compare token counts with token complexities across questions in MMLU-Pro Math, which are sorted by token complexity. The correlation between the token count of TALE-EP and the true token-complexities (Spearman $\rho=0.51$) is similar to the correlation observed for NoProperGrammar (Spearman $\rho=0.54$), which does not explicitly prompt the LLM to reason adaptively. This can help explain why both methods have similar accuracy-length tradeoffs.
As a proof of concept, we consider two prompting strategies designed to adjust reasoning effort to problem difficulty. Han et al. (2024) propose a two-step procedure called TALE-EP, which first (1) prompts the LLM to guess the minimum tokens required to solve the question and then (2) prompts the LLM to think step-by-step using the guessed number of tokens. Using our experimental results, we can compare whether adaptively adjusting the reasoning effort improves upon the accuracy-length tradeoff. The left panel of Figure 3 plots accuracy and average token-length of TALE-EP for GPT-4o on MMLU-Pro Math compared to several prompts from our experiments. Surprisingly, we observe that TALE-EP is slightly below the trade-off curve from our simple prompting strategies. More specifically, it produces a slightly worse accuracy than NoProperGrammar, even though it uses more tokens on average. This is consistent across other benchmarks (see Appendix H).
We find a potential explanation in the right panel of Figure 3, which plots the token lengths used by TALE-EP and NoProperGrammar along with the token complexity. We find that that the token counts produced by NoProperGrammar have a surprisingly high correlation with the true token complexity, which is equivalent to the correlation for the token counts of TALE-EP (Spearman $\rho=0.51$ for TALE-EP and $\rho=0.54$ for NoProperGrammar, see Appendix B for other prompts). Thus, this shows that LLMs natively adjust response length to the difficulty of the problem, even without explicitly being prompted to do so. And moreover, this capability is equivalent to more sophisticated prompting strategies, demonstrating that the simple prompting strategies we test are surprisingly strong baselines and that LLMs may struggle to estimate token complexity accurately. Nonetheless, methods based on fine-tuning (e.g. Kang et al. (2024) or TALE-PT in Han et al. (2024)) may be able to outperform these simple prompting strategies, which we leave to future research.
This result suggests that it may be challenging to achieve a better accuracy-length tradeoff. Nonetheless, if one has access to a perfect verifier (Brown et al., 2024; Setlur et al., 2024), we can substantially improve the tradeoff. Verifier Routing first (1) prompts the LLM to without chain-of-thought (NoCoT) to obtain an initial solution and (2) if the solution is incorrect it uses a longer chain-of-thought prompt (e.g. DefaultCoT) to produce a better answer. The blue dots in the left panel Figure 3 show performance of Verifier Routing with a selection of four longer prompts. This routing strategy achieves a significantly better trade-off between reasoning length and performance, and approaches the theoretical upper bound. Altogether, this shows that the upper bound $\alpha_{\pi}^{*}(T)$ can be approached through adaptive compression, although this requires a very accurate signal of problem difficulty and motivates further research.
@@ 6 Conclusion
Our study presents a systematic investigation into the trade-off between reasoning length and performance in large language models (LLMs), across different prompts. We demonstrate that this trade-off follows a universal Pareto curve, suggesting that reasoning length, rather than specific compression strategies, primarily determines accuracy. Introducing the concept of token complexity, we find that LLM performance at the question-level exhibits a sharp threshold behavior – it is successful only if its chain-of-thought is above a token threshold. Our analysis, inspired by rate-distortion theory, reveals that existing prompt-based compression strategies operate far from the optimal accuracy-length frontier, highlighting substantial room for improvement. Our work enables researchers to contextualize the performance of new methodologies for improving chain-of-thought compression and assess adaptivity of LLM reasoning.
Appendix A Limitations
@@
Appendix B Correlation with Token Complexity
The following is a table describing the Spearman correlation between token-lengths with token-complexity across different prompts for GPT-4o on MMLU-Pro Math. While the ordering of which prompts have the highest correlation changes across different models and benchmarks, we observe widely that the correlation ranges between 0 - 0.6.
@@ Appendix B Correlation with Token Complexity
Spearman $\rho$
CharLimit(50) 0.57
NoSpaces 0.56
OnlyNumbers 0.55
CharLimit(100) 0.55
NoProperGrammar 0.54
TokenLimit(10) 0.38
WordLimit(15) 0.37
WordLimit(5) 0.29
WordLimit(10) 0.28
NoCoT 0.08
Table 5: GPT-4o on MMLU-Pro Math: Spearman correlation of token-lengths with token-complexity across different prompts.
In Figure 4, we consider the average token lengths of GPT-4o on MMLU-Pro Math. We split the problems into two categories: problems that the LLM successfully solves without chain-of-thought (NoCot), and the problems NoCot unsuccessfully solves. Intuitively, the first class of problems are ‘easy’, since they do not require any chain-of-thought, and the rest are harder. Across all the prompts we consider, we observe that the average token-length among ‘easy’ problems is universally smaller than among ‘harder’ problems. Surprisingly, this is even true for the WordLimit/TokenLimit/CharLimit/etc. prompts, which are supposed to limit the LLM’s response to a fixed length. Nevertheless, despite the evidence of adaptive response lengths, this also shows that there is a lot of room for improvement: even though the LLM can solve the problem without any chain-of-thought, the model still proceeds to generate long chains-of-thought.
Figure 4: Do LLM’s tailor response length to problem difficulty? Average token length produced by GPT-4o on MMLU-Pro Math across prompts in Table 1, split by problems which can be solved without chain-of-thought and problems which NoCoT does not successfully solve. Response lengths are consistently higher for the problems NoCoT unsuccessfully solves.
Appendix C Accuracy and Average Token Length
Below are tables which describe the accuracy and average token length for several prompts across models and benchmarks.
@@ Appendix C Accuracy and Average Token Length
Appendix D Proof of Theorem 1
@@
Appendix E Example Prompt
@@
Answer the following question. PROMPT Question: QUESTION The last line of your response should be of the following format: ’Answer: ANSWER’ (without quotes) where ANSWER is your final answer.
Appendix F Tradeoff curves for more models and benchmarks